OpenStudy (anonymous):

You have a 3-card deck containing a king, a queen, and a jack. You draw a random card, then put it back and draw a second random card. Calculate the probability that you draw exactly 1 king.

OpenStudy (anonymous):

pretty small deck

OpenStudy (anonymous):

you could draw a king on the first try, and then not a king on the second that probability is \(\frac{1}{3}\times \frac{2}{3}=\frac{2}{9}\)

OpenStudy (anonymous):

haha that is what i was thinking

OpenStudy (anonymous):

or it could happen in reverse not king, then a king probability is the same since multiplication is commutative it is also \(\frac{2}{9}\)

OpenStudy (anonymous):

your final job is to add up those two numbers, aka multiply by 2

OpenStudy (anonymous):

so 4/18 but it will just give you 2/9 after reducing

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

c'mon since when is \(\frac{2}{9}+\frac{2}{9}=\frac{4}{18}=\frac{2}{9}\) ??

OpenStudy (anonymous):

must be some new math

OpenStudy (anonymous):

so it would be 4/9

OpenStudy (anonymous):

two ninths plus two ninths is four ninths just like two apples and two apples is four apples

OpenStudy (anonymous):

and i'm sorry jeesh, i was just confused..

OpenStudy (anonymous):

yeah \[\Huge \frac{4}{9}\]

OpenStudy (anonymous):

ah don't let me give you a hard time, it is late i know

OpenStudy (anonymous):

aww thanks @satellite73 just wanna thank you again for helping me

OpenStudy (anonymous):

yw got any more?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

k lets do one math in august, ick

OpenStudy (anonymous):

You flip 9 fair coins. Amazingly, the first 8 flips all come up heads. What is the probability that the final flip will be a head too?